The price of smartphone is currently rs 200 and the quantity demanded is 4m. Next year the price false to rs 180 and the quantity demanded rises to 6m. What us the price elasticity of demand?


1
Expert's answer
2021-09-24T11:29:32-0400

Price elasticity of demand (Ed)=

(Change in quantity demand/ change in price) "\times" ( price (P0) / quantity demand (Q0) )

"P_0 = 200 \\

Q_0=4 \\

P_1 = 180 \\

Q_1 = 6 \\

E_d = \frac{6-4}{180-200} \times \frac{200}{4} \\

= \frac{2}{20} \times 50 \\

= 5"